OpenStudy (anonymous):

Matt drove at a steady speed on the first day of his vacation. The table shows data about the first day. When matt stopped for lunch, he had traveled 232 miles. How long did he drive before lunch?

OpenStudy (anonymous):



OpenStudy (compassionate):

I'm going to help you solve this, okay?

OpenStudy (anonymous):

alright cool

OpenStudy (compassionate):

We know his time and speed, we need to find his distance(d) Start off by finding his time using the formula: d = r*t ----------- d = 1.5*87, d = 130.5 ----------- d = 2.25*130.5 d = 293.7 ----------- d = 3*174 d 522 _____________ So we know the distance for his 1, 2, and 3rd log.

OpenStudy (compassionate):

Can you take it from here?

OpenStudy (compassionate):

Also, is there any additional information you're leaving out?

OpenStudy (compassionate):

Pardon me, I meant to do the rate formula.

OpenStudy (compassionate):

You want to divide distance by time and get the rate you traveled.